OpenStudy (anonymous):

Solve the equation: 2x + 7 = (x + 3)4

OpenStudy (anonymous):

\[2x+7=4x+12\] \[7=2x+12\] \[-5=2x\] \[x=-\frac{5}{2}\]

OpenStudy (anonymous):

Question Stalker <3

hero (hero):

Satellite, did Jim Thompson give you his program?

OpenStudy (anonymous):

i'm on a roll

OpenStudy (anonymous):

program?

OpenStudy (anonymous):

@hero chat?